A 15 inch television set has a width of 12 inches. What is its' height?
Thank you!

Answer by rfer(16322)

a^2=C^2-b^2
a^2=15^2-12^2
a^2=225-144
a^2=81
a=9 in
TV's are measured on the diaginal.